OpenStudy (anonymous):

x/5 - y/6 = 1 can I graph this?

OpenStudy (anonymous):

Yes, It is a straight line. What I recommend is getting rid of the fractions. So multiply the whole equation by 30

OpenStudy (anonymous):

so my first point would be 5 on the x axis and then 6 on the y axis

jimthompson5910 (jim_thompson5910):

x/5 - y/6 = 1 x/5 + y/(-6) = 1 The equation is now in the form x/a + y/b = 1 where 'a' is the x-intercept and 'b' is the y-intercept. So you know that the two points (5, 0) and (0, -6) are on this line. This is because a = 5 and b = -6.

OpenStudy (anonymous):

k got that but I get confused when graphing it

jimthompson5910 (jim_thompson5910):

so plot (5,0) and (0,-6) to get